. If you are running a t-test for dependent means on a group of 25 individuals, your degrees of freedom will be which of the following? a) 25 b) 26 c) 24 d) 12.5
Answer #1

If you are running a t-test for dependent means on a group of 25 individuals.

Sample size = n = 25

Degrees of freedom = n - 1 = 25 - 1 = 24

c) 24